OpenStudy (anonymous):

tan9 - tan27 - tan63 + tan 81 = ?

OpenStudy (lgbasallote):

hmmm they're all factors of 9....

hartnn (hartnn):

NOTE : 81+9=90 63+27=90

OpenStudy (lgbasallote):

ahh yes..that...

OpenStudy (anonymous):

that wont help...@hartnn i think...

hartnn (hartnn):

really ? tan 81 = cot 9 = 1/ tan 9 now??

OpenStudy (anonymous):

lol....Go on..u will.....see....

hartnn (hartnn):

hmmm.....sin 18 and sin 54...

OpenStudy (anonymous):

The answer should be 4

hartnn (hartnn):

u need sin 18 = \(\sqrt5-1\) else u cannot solve it.

OpenStudy (anonymous):

Cant we Do..this... With out Knwing sin 18....i mean,,,,any other method

OpenStudy (anonymous):

tan 9=sin9/cos9=cos81/sin81=cot81 tan 27=sin27/cos27=cos63/sin63=cot27 so! tan9+cot81-tan63-cot27=? remember the relation b/n tan and cot of complementary angles tan a=cot 90-a

OpenStudy (anonymous):

\[\tan9 - \tan 27 - \tan63 - \tan81 \] \[ \tan9 + \tan81 - \tan27 - \tan63\] \[\frac{\sin9}{\cos9}+\frac{\sin 81}{\cos 81}-\frac{\sin 27}{\cos27}-\frac{\sin63}{\cos63}\] \[ \frac{\sin90}{ \cos81\cos9} -\frac {\sin90}{ \cos63\cos27}\] \[ \frac{1}{\sin9\cos9} - \frac{1}{\sin27\cos27}\] \[ \frac{2}{\sin18} - \frac{2}{\sin54}\] \[ 2\frac{\sin54 - \sin18}{\sin18\sin54}\] \[ \frac{4\cos36\sin18} {\sin18\cos36}=4\]